OpenStudy (anonymous):

Find the volume of the solid generated by revolving the region bound by the graphs of the equations around the line x=5. y=5-x, y=0, y=4, x=0

OpenStudy (amistre64):



OpenStudy (amistre64):

id shell it

OpenStudy (anonymous):

What integral do I take for using discs though?

OpenStudy (amistre64):

disc is just using circles instead of cylindars;\[\int pi f(x)^2dx\]

OpenStudy (amistre64):

but it might be best to set it up better, our radius is in the form \(f^{-1}(x)\) and there is an inner and outer radius to account for

OpenStudy (amistre64):

y=5-x x=5-y x-5=-y x+5=y would be our radius function: or y+5=x might be a better notation

OpenStudy (amistre64):

the outer radius is just "5" the inner radius is what we have to remove, right?

OpenStudy (amistre64):